### v1

# VulMCI : Code Splicing-based Pixel-row Oversampling for More Continuous Vulnerability Image Generation

Abstract.
In recent years, the rapid development of deep learning technology has brought new prospects to the field of vulnerability detection. Many vulnerability detection methods involve converting source code into images for detection, yet they often overlook the quality of the generated images. Due to the fact that vulnerability images lack clear and continuous contours, unlike images used in object detection, Convolutional Neural Networks (CNNs) tend to lose semantic information during the convolution and pooling processes. Therefore, this paper proposes a pixel row oversampling method based on code line concatenation to generate more continuous code features, addressing the issue of discontinuity in code image coloration.Building upon these contributions, we propose the vulnerability detection system VulMCI and conduct tests on the SARD and NVD datasets. Experimental results demonstrate that VulMCI outperforms seven state-of-the-art vulnerability detectors (namely Checkmarx, FlawFinder, RATS, VulDeePecker, SySeVR, VulCNN, and Devign). Compared to other image-based methods, VulMCI shows improvements in various metrics, including a 2.877% increase in True Positive Rate (TPR), a 5.446% increase in True Negative Rate (TNR), and a 5.91% increase in Accuracy (ACC). On the NVD real-world dataset, VulMCI achieves an average accuracy of 5.162%, confirming its value in practical vulnerability detection applications. We will share the source code of the paper on GitHub after it is accepted.
1. Introduction
With the widespread use of computer networks and the ubiquity of the Internet, vulnerability detection has become increasingly crucial as malicious hackers and cybercriminals continuously seek novel methods to infiltrate systems and pilfer sensitive information. Furthermore, as software scales and complexities continue to grow, vulnerabilities within source code have become more insidious and pervasive, posing a severe threat to information security. According to the latest data from the National Vulnerability Database (NVD)(of Standards and Technology, 2023), we can observe that in 2021, the number of disclosed vulnerabilities surpassed the twenty-thousand mark, and even though the figure decreased to slightly over thirteen thousand in 2022, it remains a substantial quantity. Consequently, the need for efficient and automated source code vulnerability detection has emerged as an urgent requirement to ensure the robustness and reliability of software systems.
Traditional vulnerability detection methods typically employ techniques based on code similarity(Jang et al., 2012; Kim et al., 2017; Li and Ernst, 2012; Li et al., 2016; Pham et al., 2010) or rule-based approaches(FlawFinder, 2021; Rough Audit Tool for Security, 2021; Checkmarx, 2021) to analyze source code in order to identify potential vulnerabilities. However, these conventional methods exhibit several notable limitations. Firstly, they often require a significant amount of manual work and rule formulation, which becomes impractical when dealing with large-scale code repositories. Secondly, they frequently struggle to address complex vulnerability types, especially those with concealment and variants. Finally, these methods tend to produce false positives and false negatives when handling extensive code repositories, thus compromising the accuracy and efficiency of the detection process.
Nevertheless, in recent years, the rapid advancement of deep learning technology has brought new prospects to the field of vulnerability detection. Deep learning methods have the ability to automatically learn patterns and features from extensive source code without the need for manual intervention, rendering them highly scalable. Furthermore, deep learning technology is effective in handling complex vulnerability types, thereby enhancing the accuracy of vulnerability detection. These advantages have positioned deep learning-based vulnerability detection methods as a current focal point of research and attention.
Existing deep learning-based vulnerability detection systems utilize various methods, including processing code into textual representations using natural language processing(Li et al., 2018; Russell et al., 2018b), analyzing code structure graphs using graph neural networks(Duan et al., 2019a; Zhou et al., 2019), and generating images from code for detection(Wu et al., 2022). While each of these methods has its merits, there is still room for improvement. In some prior studies(Li et al., 2018, 2022), custom Common Weakness Enumeration (CWE) vulnerability sets were used to match and locate vulnerabilities. However, this approach may lead to incomplete vulnerability coverage and instances where vulnerability labels do not match in real-world datasets. Additionally, many methods(Li et al., 2022; Lv et al., 2023) rely on the concept of code slicing to eliminate redundant information, depending not only on highly accurate vulnerability line localization but also risking the loss of complete program semantic information during slicing.
Furthermore, methods that transform source code into images, such as VulCNN(Wu et al., 2022), although preserving the complete semantics and structural information of functions, have been found through experiments to generate RGB images with significant instances of black and discontinuous scattered color points in the channels. This characteristic is considered detrimental to the classification task of Convolutional Neural Network (CNN) models.
In this paper, we adopt the baseline method of converting source code into images and make improvements to address the issue of discontinuity in image coloration. It is important to note that we do not simply apply morphological operations to the images, as doing so would disturb the weight differences between sentences. Instead, we apply a pixel row oversampling algorithm to source code lines with dependencies to enhance the continuity between them. The main advantage of this approach is the reduction of data noise, thereby minimizing differences between adjacent data points. This results in a more stable model that is less sensitive to small variations in the input data, contributing to improved model stability and generalization capabilities.
In summary, this paper makes the following contributions:
We identified a novel issue wherein there are discontinuous and abrupt numerical changes when code is transformed into image representations. This discontinuity can potentially have a negative impact on the training of Convolutional Neural Network (CNN) models, particularly when a significant amount of line information is lost after the image undergoes pooling operations.
We propose a pixel row oversampling algorithm based on code concatenation. To address the issue of discontinuous color points in images directly generated from code, we employ an adaptive segmentation method to insert concatenated rows with weights between adjacent rows. This approach balances local variations in the image, highlighting important features and enhancing the CNN’s ability to extract key features, ultimately contributing to improved classification performance.
We introduce a new system, VulMCI, which can extract function samples from complex source code and generate code feature images with more continuous upper and lower pixel rows.
We introduce a new hypothesis that high-quality vulnerability detection images can enhance the detection performance of Convolutional Neural Network (CNN).
Paper organization. The remaining sections of this paper are organized as follows. Section 2 introduces our method. Section 3 presents the experimental results. Section 4 discusses future work. Section 5 provides an overview of related work. Finally, Section 6 summarizes the key findings of this paper.
2. Our Method
The existing methods for generating images from code overlook the continuity of pixel rows, resulting in suboptimal image quality. Therefore, we propose a systematic approach in VulMCI to generate high-quality images with more continuous upper and lower pixel rows. As illustrated in Figure1, VulMCI comprises four main stages: building function samples, pixel-row oversampling, code encoding and embedding, and generating three-channel images.
Building function samples: Our system supports the extraction of function-level slice samples from initial program files, which undergo normalization. This process involves three main steps: removing comments, standardizing function names, and standardizing variable names.
Pixel-row oversampling: In existing functions, we employ an adaptive segmentation method to insert new concatenated rows between adjacent rows.
Generating three-channel images: We adopt three centrality analysis indicators from VulCNN(Wu et al., 2022) (degree centrality(Freeman, 1978), katz centrality(Katz, [n. d.]), closeness centrality(Freeman, 1978)). Different from VulCNN, which generates Program Dependency Graph (PDG), we generate Code Property Graph (CPG) for functions, additionally incorporating Abstract Syntax Tree (AST) and Control Dependency Graph (CDG).
2.1. Pixel-row oversampling
Given adjacent code lines, denoted as Line and Line (where represents the line number), adaptive concatenation is performed between these two adjacent lines. The concatenation process is carried out as follows:
Step 1: In Line , take the number obtained from Formula 1 as the length of the substring, and obtain the substring .
(1) -
Step 2: In Line , take the number obtained from Formula 2 as the length of the substring, and obtain the substring .
(2) -
Step 3: Obtain the new concatenated line , and complete the words at the character-level separation points.
For , k-1 new concatenated lines will be gradually inserted between adjacent lines. For example, when , the values for are 1 and 2, and two concatenated lines will be inserted between each adjacent pair of lines.
In the red-boxed portion of the function samples shown in Figure 1, where , (including the number of spaces between words), , , and , , thus only one concatenated line will be inserted. For ,= and &=.
, . The resulting concatenated line is .
2.2. Generating a three-channel image
We inputted function samples into the Joern tool(Yamaguchi et al., 2014), which utilizes static analysis to generate a Code Property Graph (CPG). The CPG includes Control Flow Graph (CFG), Control Dependency Graph (CDG), Data Dependency Graph (DDG), and Abstract Syntax Tree (AST). Similar to VulCNN, we derived an adjacency matrix and calculated in-degrees, out-degrees, and degrees for all code lines based on the edge and node information in CPG. Subsequently, we conducted centrality analysis on the CPG, computing the degree centrality, Katz centrality, and closeness centrality for each code line node. The formulas for these three centrality metrics are as follows, following similar procedures as in VulCNN.
Degree Centrality(Freeman, 1978):
(3) Where is the degree of node , and is the number of nodes in the graph.
Katz Centrality(Katz, [n. d.]):
(4) Where and are constants, is the adjacency matrix of the graph, and is the centrality of the -th node.
(5) Where is the largest eigenvalue of the adjacency matrix .
Closeness Centrality(Freeman, 1978):
(6) Where is the distance between nodes and , and is the number of nodes in the graph.
The formula for the centrality metrics of the newly generated concatenated line is as follows:
2.3. algorithmic descriptions
Algorithm 1 outlines how VulMCI transforms function code into a three-channel image. Firstly, we preprocess the function code, and then we perform adaptive segmentation based on to concatenate and insert new lines between adjacent code lines, forming new lines. Subsequently, we calculate centrality metrics for each code line based on the Code Property Graph (CPG). Following this, each of these code lines is vectorized using sent2vec, producing corresponding vector representations. These vectors are then multiplied by three expanded lists of centrality metrics and mapped into three pixel channels resembling RGB. Ultimately, this process creates a code image with improved continuity.
2.4. Theoretical Analysis
Definition 2.0 ().
For any two adjacent pixel rows represented by vectors and , an image is considered pixel-row continuous when the absolute difference between corresponding elements, denoted as , where gap is a small value.
Theorem 2.2 ().
Our method has a good probability of producing more continuous images.
Proof.
Step 1: We demonstrate the continuity between adjacent pixel rows and concatenated pixel rows under specific conditions. The objective of the sent2vec(Pagliardini et al., 2018) model is to learn word and sentence embedding by minimizing a loss function. The simplified representation of the objective is as follows:
where represents the word embedding matrix, represents the sentence embedding matrix, is the corpus, and represents the word index list corresponding to the sentence in the corpus.The parameterized representation of the sentence vector is as follows. We aim to obtain a set of parameters to demonstrate the effectiveness of our concatenated rows.
If for any pair of elements, and holds, then the two vectors and the concatenated vector are considered continuous.
Step 2:We establish that the UV parameters in the training objective of the sent2vec model satisfy the requirements of the above inequality.
The complete training objective of the Sent2Vec(Pagliardini et al., 2018) model, an unsupervised learning method for learning and inferring sentence embeddings, is as follows:
where and represent the embedding matrices for words and sentences respectively, is the corpus containing the set of sentences to be learned. Additionally, represents a specific sentence in the corpus, and is the target word in sentence . represents the embedding vector for the target word, and is the embedding vector for the sentence with the target word removed. represents the set of negative samples extracted for the target word . The function is the binary logistic regression loss function, measuring the model’s predictive performance on positive and negative samples. By minimizing this loss function, the model learns embeddings that effectively capture semantic relationships between words and sentences in an unsupervised manner.
The first part of the loss function focuses on positive samples, encouraging similar sentences to be closer in the embedding space through the logistic regression loss function . The second part addresses negative samples, pushing dissimilar sentences further apart in the embedding space through the logistic regression loss function . The overall objective is to minimize the loss function by learning and , resulting in embeddings with enhanced semantic relationships between sentences and words. By concatenating adjacent code tokens and inserting them in between, more context information is introduced, making adjacent code lines closer in the embedding space. The encoding and embedding stage in Figure 1 illustrates this characteristic. Thus, the training objective of sent2vec yields a set of UV parameters, theoretically resulting in a small gap. This demonstrates the continuity between adjacent vector rows in our approach. ∎
2.5. Classification
In the realm of vulnerability detection, the adoption of advanced techniques becomes imperative to handle the complexity and nuances of source code. Convolutional Neural Networks (CNN) have emerged as a powerful tool in various image-based tasks, showcasing their ability to discern intricate features and patterns. In this section, we leverage the capabilities of CNN to address the challenges inherent in source code vulnerability detection.
After completing the image generation phase, we transformed the source code of functions into images. For a given image, we initiated training of a CNN model, subsequently applying this model for vulnerability detection. Since CNN models require input images of uniform size, and the number of code lines in different functions may vary, it becomes essential to select a suitable threshold length for extension or cropping.
The code length distribution of function samples in the experimental dataset (SARD) is illustrated in Figure 2. It is observed that the majority of function lengths are less than 100 lines. Therefore, we have chosen a code length threshold of 100 lines to generate our input images. And previous experiments have shown that choosing 100 lines as a threshold can balance overhead and detection performance well. Furthermore, we dynamically adjust the parameter for the concatenated lines, setting the code line threshold to .We take 100 lines as an example. For functions with code lengths less than 100, we pad the vectors with zeros at the end. In the case of functions exceeding 100 lines, we truncate the trailing portion of the vectors. The input image size is set to 3×100×128, where 3 corresponds to three channels (i.e., ”degree channel,” ”Katz channel,” and ”proximity channel”), 100 is the code line threshold, and 128 represents the dimensionality of the sentence vectors.
After generating images of fixed size, we build a Convolutional Neural Network (CNN) model through training on these images. As depicted in Figure 3, we employ convolutional filters of varying shapes, with dimensions m * 128, to ensure each filter can extract features across the entire space of the embedded sentences. The size of the filters roughly determines the length of the sentence sequences considered simultaneously. In VulMCI, we select 10 filters of different sizes (from 1 to 10), each containing 32 feature maps to capture features from different parts of the image. Following the max-pooling operation, the length of our fully connected layer is 320. Table 1 provides a detailed description of the parameters used in VulMCI. The entire model adopts Rectified Linear Unit (ReLU)(Dahl et al., 2013) as the non-linear activation function. Additionally, we utilize cross-entropy loss as the loss function in CNN for penalizing incorrect classifications. We employ Adam(Kingma and Ba, 2014) as the optimizer with a learning rate set to 0.001. Once the training is completed, we use the trained CNN model to classify new functions, determining whether they possess vulnerabilities.
| parameters | settings |
| loss function | Cross Entropy Loss |
| activation function | ReLU |
| optimizer | Adam |
| batch size | 32 |
| learning rate | 0.001 |
| epoch num | 100 |
3. EXPERIMENTS
In this section, our aim is to address the following research questions:
RQ1: How much does the detection performance improve with different values of the parameter in the pixel-row oversampling operation?
RQ2: How does the detection performance of VulMCI compare to other state-of-the-art vulnerability detection systems?
RQ3: Can VulMCI be applied to real-world vulnerability scanning?
3.1. Experiment Settings
The dataset used in this paper is sourced from the vulnerability dataset provided by SySeVR(Li et al., 2022). This dataset originates from SARD(sar, 2018) and NVD(of Standards and Technology, 2023), encompassing a total of 126 vulnerability types, each uniquely identified by a Common Weakness Enumeration (CWE) ID(cwe, 2018). The SARD dataset comprises production, synthetic, and academic programs (referred to as test cases) categorized as ”good” (i.e., without vulnerabilities), ”bad” (i.e., containing vulnerabilities), and ”mixed” (i.e., vulnerabilities with available patch versions). The NVD dataset includes vulnerabilities in 19 popular C/C++ open-source products (software systems), along with possible diff files describing the differences between susceptible code and its patched versions. The SARD dataset contains a total of 21,233 program files, while the NVD dataset includes 2,011 real vulnerabilities and their corresponding fixed version files. Our system processed the dataset by extracting function samples from program files and removing function samples with less than ten lines of code. This exclusion was based on the observation that the majority of functions with fewer than ten lines only involve external function calls and cannot be easily classified as usable function samples. Consequently, our final processing resulted in 13,165 samples of vulnerable functions and 5,612 samples of non-vulnerable functions.
In this study, we employed a k-fold cross-validation (k=5) strategy to partition the dataset, creating five mutually exclusive subsets. Four subsets were used for training the model, while one subset was reserved for testing. This partitioning approach ensures the utilization of distinct data subsets for both training and testing phases, facilitating the assessment of the model’s generalization performance. The devices and experimental parameters employed in this study are presented in Tables 1 and 2, respectively. To comprehensively evaluate the performance of our proposed method, we utilized multiple evaluation metrics, including False Positive Rate (FPR), False Negative Rate (FNR), Precision (Pr), Recall (Re), F1 Score (F1), and Accuracy (ACC). Each metric provides detailed insights into the algorithm’s performance across different aspects, offering a nuanced understanding of its strengths and limitations.
| Device | Type | Version |
| GPU | Tesla | V100 |
| CPU | Intel Xeon Silver | 6130 |
| Operating system | Centos linux release | 7.9.2009 |
| Package | Python | 3.8 |
| Torch | 1.12.1 | |
| Matplotlib | 3.8.2 | |
| Numpy | 1.25.2 | |
| Networkx | 3.1 | |
| Joern | 2.0.121 |
3.2. Experimental Results
3.2.1. Experiments for Answering RQ1
VulMCI, the method proposed by us for vulnerability detection, demonstrates excellent performance across multiple evaluation metrics, achieving an F1 Score of 92.645% and an accuracy (ACC) of 89.157%. Figure 4 illustrates the changes in loss value and accuracy during the training of VulMCI over 200 epochs. It can be observed that the loss value gradually decreases and levels off around 0.11, while the training accuracy steadily increases and fluctuates around 92.8%. This provides evidence of the effectiveness of our model. To assess the impact of the concatenated line parameter , we conducted a parameter analysis experiment, and the results are presented in Table 3.
| k | FPR | FNR | Pr | Re | F1 | ACC |
| 1 | 21.429 | 9.23 | 92.221 | 90.77 | 91.49 | 87.559 |
| 2 | 20.476 | 8.677 | 92.583 | 91.323 | 91.949 | 88.217 |
| 3 | 20.595 | 8.464 | 92.559 | 91.536 | 92.044 | 88.342 |
| 4 | 17.738 | 9.315 | 93.468 | 90.685 | 92.055 | 88.468 |
| 5 | 18.69 | 8.847 | 93.174 | 91.153 | 92.152 | 88.562 |
| 6 | 20.476 | 8.592 | 92.589 | 91.408 | 91.995 | 88.28 |
| 7 | 19.048 | 8.805 | 93.056 | 91.195 | 92.116 | 88.499 |
We evaluated the effectiveness of concatenated lines by setting 7 values for in the analysis. As shown in Table 3, when ranges from 1 to 5, larger values of indicate more inserted concatenated lines, resulting in enhanced image continuity and gradually improving detection accuracy. Compared to (no inserted concatenated lines) with an accuracy (ACC) of 87.559%, the concatenation operation brings about a roughly 1% improvement. When , the detection performance fluctuates but overall remains at a high level.
3.2.2. Experiments for Answering RQ2
In this section, we compare VulMCI with several vulnerability detection tools, including a commercial static vulnerability detection tool (i.e., Checkmarx(Checkmarx, 2021)), two open-source static analysis tools (i.e., FlawFinde(FlawFinder, 2021) and RATS(Rough Audit Tool for Security, 2021)), and four deep learning-based vulnerability detection methods (i.e., VulDeePecker(Li et al., 2018), VulCNN(Wu et al., 2022), Devign(Zhou et al., 2019), and SySeVR(Li et al., 2022)). Figure 5 presents a bar chart that intuitively illustrates the metrics of True Positive Rate (TPR), True Negative Rate (TNR), and Accuracy (ACC) for each of these methods.
For the three previous static analysis tools, RATS, Checkmarx, and FlawFinder, their detection performance is not satisfactory, with accuracy all below 60%. One possible explanation is that they rely on human experts to define vulnerability rules, and as vulnerability types become more complex, human experts may not be able to comprehensively define all patterns, leading to higher false positives and false negatives.
For the VulDeePecker and SySeVR methods, both employ program slicing on the dataset. They vectorize the slices and use them to train Bidirectional Recurrent Neural Networks (BRNN) for vulnerability detection. The performance disparity between these two systems stems from VulDeePecker solely considering vulnerabilities caused by API function calls and only accounting for data dependency information. On the other hand, SySeVR incorporates semantic information caused by both data and control dependencies. It supports a broader range of vulnerability syntax labels for matching vulnerabilities. Consequently, SySeVR achieves a 9.47% higher accuracy than VulDeePecker. SySeVR performs well on the experimental dataset, partly because its set of Common Weakness Enumeration (CWE) vulnerability syntax labels covers most vulnerabilities in the dataset. However, its effectiveness in detecting unknown or real vulnerabilities might be limited. In other words, the coverage of the vulnerability label set is restricted, and vulnerabilities might span multiple lines, making precise localization and slicing challenging. In this experiment, the accuracy is 2.064% lower compared to VulMCI.
Devign utilizes graph neural networks, leveraging node relationships in the graph structure. However, due to its use of complex network node information in the compound graph, it introduces redundant information weakly related to vulnerability nodes, increasing model complexity. Experimental results also show that VulMCI outperforms Devign across various metrics, with an 8.315% increase in True Negative Rate (TNR) and a 7.415% improvement in accuracy. While Devign can identify new vulnerabilities to some extent, it may face challenges when handling large-scale code repositories. VulCNN, using centrality analysis, transforms time-consuming graph analysis into efficient image scanning. It effectively addresses large-scale scanning, but ignoring the discontinuity of pixel information in images results in insufficient vulnerability feature extraction. VulMCI retains richer code relationships in the Code Property Graph (CPG), leading to a 2.877% improvement in True Positive Rate (TPR), a 5.91% increase in TNR, and a 5.162% improvement in accuracy.
While Devign can identify new vulnerabilities to some extent, it may face challenges in handling large-scale code repositories. VulCNN transforms time-consuming graph analysis into efficient image scanning through centrality analysis. However, due to redundant information in the generated image and the discontinuity of pixel information, VulMCI improves upon VulCNN’s weaknesses. VulMCI achieves a 19.601% increase in TPR, a 5.446% increase in TNR, and a 9.231% increase in accuracy, while absorbing its advantages.
3.2.3. Experiments for Answering RQ3
To assess the real-world vulnerability detection capability of VulMCI, we conducted model training and testing on 2011 real vulnerabilities and their corresponding patched versions from the National Vulnerability Database (NVD), maintaining the same data partitioning strategy as in previous experiments. Due to the scarcity of real-world vulnerability data, and to mitigate the potential impact of individual sample variations on result accuracy, we performed five repetitions of the experiments. We saved the results of the final generation of the model, conducted testing on each fold of the test set, calculated the average for each fold, and ultimately obtained the results from five model testing sessions.
To precisely evaluate the model’s generalization ability, we computed the average accuracy and accuracy standard deviation for each experiment. The stability of the experimental results was visually presented using an error bar graph, as depicted in Figure 6. The error bar graph vividly illustrates the average accuracy of each experiment and expresses the dispersion of accuracy in the form of error bars. The length of the error bars provides insights into the fluctuation in model performance across different experiments. Shorter error bars indicate relatively stable model performance, while longer error bars signify higher instability in performance.
The experimental outcomes reveal that VulMCI exhibits strong detection performance on real-world datasets. The average accuracy across five repeated experiments reached 86.88%, with minimal error fluctuations. The model’s performance varied within the range of 0.77% to 0.83%, indicating its adaptability to diverse data distributions. These findings not only offer robust support for practical applications but also suggest promising prospects for real-world deployment of VulMCI.
4. Discussion
The DISCUSSION section highlights key aspects of the entire system process and outlines future directions for development:
4.1. Limitations in Static Analysis Efficiency
Throughout the entire system workflow, the computational cost of static analysis for generating the Code Property Graph (CPG) is a major bottleneck, requiring over twenty hours for a dataset with twenty thousand function samples. We are currently contemplating the reconstruction of a new, lightweight code analysis tool to enhance operational efficiency.
4.2. Pixel Row Oversampling Focus
While enhancing code line continuity, we currently overlook pixel row internal continuity. Ongoing research aims to validate the effectiveness of this concept.
4.3. Future Expansion and Scalability
In subsequent phases, we plan to extend our method to other graph-based detection systems to further validate its scalability. This extension aims to assess the adaptability and effectiveness of our approach beyond the current system.
5. RELATED WORK
In recent years, researchers have proposed various methods for vulnerability detection, categorizing static detection techniques into code similarity-based and pattern-based methods. Code similarity-based approaches, such as (Kim et al., 2017; Li and Ernst, 2012; Pham et al., 2010; Sajnani et al., 2016), are generally suitable for identifying code clones or repetitions but exhibit poor performance in detecting unknown vulnerabilities.
Pattern-based methods encompass several directions. Rule-based methods rely on human experts to manually construct vulnerability features. Tools like Checkmarx(Checkmarx, 2021), FlawFinder(FlawFinder, 2021), and RATS(Rough Audit Tool for Security, 2021) using this approach often struggle to ensure coverage in practical detection, requiring a high level of expertise and thus yielding suboptimal results. Early machine-learning-based vulnerability detection methods, such as those based on metric program representations, typically employed feature engineering to manually extract vulnerability-related features. Classic methods include code churn(Perl et al., 2015; Zimmermann et al., 2010; Morrison et al., 2015; Meneely et al., 2013; Shin et al., 2011; Shin and Williams, 2013; Walden et al., 2014), code complexity(Zimmermann et al., 2010; Moshtari et al., 2013; Morrison et al., 2015; Younis et al., 2016; Shin et al., 2011; Shin and Williams, 2013; Walden et al., 2014), coverage(Zimmermann et al., 2010; Morrison et al., 2015), dependency(Zimmermann et al., 2010; Morrison et al., 2015), organizational(Zimmermann et al., 2010; Morrison et al., 2015), and developer activity(Perl et al., 2015; Bosu et al., 2014; Meneely et al., 2013; Shin et al., 2011; Meneely and Williams, 2010). In recent years, researchers have started using deep learning methods to automatically extract vulnerability features.
Due to different program representations at various compilation stages, deep learning-based program representation methods can be categorized into three types based on the organizational form of code representation: sequence-based(Grieco et al., 2016; Wu et al., 2017; Russell et al., 2018a; Yan et al., 2021; Li et al., 2021a; Tian et al., 2020; Li et al., 2021c), syntax tree-based(Li et al., 2019; Tanwar et al., 2021; Zhang et al., 2021; Wang et al., 2016; Lin et al., 2017), and graph-based(Duan et al., 2019b, 2020; Cao et al., 2020; Li et al., 2021b; Cheng et al., 2021; Zheng et al., 2021; Wang et al., 2020; Okun et al., 2013). VulDeePecker(Li et al., 2018) slices the program, collects code snippets, and transforms them into corresponding vector representations. Ultimately, these vectors are used to train a Bidirectional Long Short-Term Memory (BLSTM) model for vulnerability detection. SySeVR(Li et al., 2022) first performs syntactic analysis and then semantic slicing on vulnerable code, transforming sliced code into fixed-length vectors fed into a Bidirectional Gated Recurrent Unit (BGRU) model for training. VulDeeLocator(Li et al., 2021c) utilizes intermediate code to define program slices for vulnerability detection, proposing a new variant of Bidirectional Recurrent Neural Network (BRNN), namely BRNN-vdl, for vulnerability detection and localization. Devign(Zhou et al., 2019) is a graph neural network-based model that encodes the original function code into a joint graph structure containing rich program semantics. The gated graph recurrent layer learns node features by aggregating and propagating information in the graph, and the Conv module extracts meaningful node representations for graph-level predictions. VulCNN(Wu et al., 2022) is a scalable graph-based vulnerability detection system that effectively converts function source code into images while preserving program semantics, suitable for large-scale vulnerability scanning.
6. CONCLUSION
In this paper, we propose and validate the effectiveness of high-quality images with strong continuity in improving the detection accuracy of Convolutional Neural Networks (CNNs). Introducing the VulMCI system method as an improvement over the baseline approach of transforming source code into images, we address the issue of color discontinuity. Within VulMCI, an adaptive stitching pixel-row oversampling method is designed to generate more continuous code features. Theoretical analysis and parameter studies are conducted, showcasing high detection performance on the SARD and NVD datasets. This multidimensional approach substantiates the validity of the experimental results.
Acknowledgment
This work was supported by the National Natural Science Foundation of China under Grant 62203337. The authors would like to express their gratitude to the creators of the SySeVR dataset and the open-source code provided by the VulCNN authors.